# natural language processing; greenwashing detection; corporate climate-related greenwashing; climate pretrained models; climate topic detection; green claim detection; climate QA; climate-related deceptive techniques; enviromental performance prediction copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language processing

### v1

@@ 6.1. Evaluation Methodology
Many tasks in greenwashing detection are framed as classification tasks. However, in some cases, the labels are not sufficiently distinguishable. For example, specificity classification (Bingler et al., 2023) is the task of distinguishing specific claims from unspecific claims. It turns out that this task is itself too unspecific, as the low inter-annotator agreement shows. However, even for tasks that seem easy to distinguish, some edge cases remain ambiguous, e.g., the distinction between “climate-related” and “not climate-related”. The climate detection model climateBERT (Bingler et al., 2023) classifies the sentence “This year, we are producing more than 10,000 [bicycle/electric bicycle], demonstrating our continued growth and commitment to meeting market demand.” as “climate-related” for bicycle but not for electric bicycle. While electric bicycles are often associated with solutions to reduce emissions, the statement is not explicitly making the link. If we consider indirect links, then non-electric bicycles are also solutions. There lies the question of which product is inherently linked to climate-change, and which one requires contextualization.
When labels in classification tasks are not clearly distinguishable, the task is not well-defined. Consequently, it does not mean much if models perform well on such a task. It merely means that the model has learned whatever label distribution the dataset had. Work in this area should thus always first check human inter-annotator agreement on the task to make sure it is meaningful.
@@
Some tasks are not classification tasks, but text generation tasks. Evaluating generated text is a challenging task (Sai et al., 2022), and different techniques are being used: BLEU, METEOR, human-based evaluation, and LLM-based evaluation (Mullappilly et al., 2023; Rony et al., 2022; Vaghefi et al., 2023; Thulke et al., 2024). BLEU and METEOR are not well-suited for evaluating tasks such as question answering  (Chen et al., 2019), as they rely only on n-gram similarity. Human evaluation also has its shortcomings. First, it is limited to smaller sizes and variations in settings. Second, human evaluation of generated texts should be based on explicit guidelines (Thulke et al., 2024; Bulian et al., 2024), but sometimes resorts to comparisons (Mullappilly et al., 2023; Thulke et al., 2024) or simple user preferences (Mullappilly et al., 2023). Third, human evaluation can be influenced by factors other than the quality of the answer – for example, the observer-expectancy effect (Rosenthal, 1976). For instance, in ChatClimate (Vaghefi et al., 2023), humans evaluate the accuracy of text generated by a generic ChatGPT vs. the proposed ChatClimate chatbot. The text produced by the ChatClimate bot always gives specific references (down to page numbers), while the generic ChatGPT does not. This allows evaluators to easily identify the more precise answer based solely on the type of references used, raising the concern that this criterion may have disproportionately influenced assessments, potentially overlooking broader aspects such as overall accuracy or quality. LLM-based evaluation has its own challenges: Language models are known to reproduce human biases (Gallegos et al., 2024). They are sensitive to the way questions are formulated, to the order of answers in multiple-choice questions (Pezeshkpour and Hruschka, 2024), and to cues that can influence human evaluation such as using “Our Model”/“Their model” (which is the wording used in Mullappilly et al. (2023)). LLM-based evaluation should thus always adhere to the same rigorous standards that are applied to human evaluation.
Figure 1. Upset plot (Lex et al., 2014) of the methods used in the studies mentioned in this literature review. Each bar counts the number of studies reporting a given set of methods. The number of a papers using a given method is reported on the horizontal bars. We have highlighted in red the intersection sets that include both fine-tuned Transformers and keyword-based methods.
@@
Figure 1 shows the approaches that are commonly used in greenwashing detection tasks (more details in Table LABEL:tab:models). We see that the most common solution is fine-tuning transformer models; more recent approaches include using LLMs as few-shot learners. The figure also shows which baselines have been used:
@@
  • Random/Majority Baseline: Any proposed approach should beat a random baseline. However, we find that only 9% of papers we studied compare their results to the random or majority baseline. For example, Friederich et al. (2021) reported performances under 50% for a binary classification task, which is worse than a theoretical random classifier on a balanced dataset. Their dataset is not balanced, but it is difficult to draw conclusions without the performance of a random classifier. Random classifiers cannot just suggest that a model underperforms, but also that a model has a comparatively good performance. For example, on SciDCC or ClimaTOPIC, the performance of models is around 50-60%. While this looks mediocre, it is in fact an excellent performance as the random classifier performed around 5-7% (Spokoyny et al., 2023). In case of a heavily imbalanced dataset, the majority baseline is more meaningful. If such a baseline is not provided, but the dataset is available, it can even be computed a posteriori. Unfortunately, neither is the case for the work of Friederich et al. (2021). Similarly, a constant value baseline can help understand the performance of a regression model. Sharma et al. (2024) and Bronzini et al. (2023) evaluated the prediction of the ESG score, but do not provide such a baseline. Thus, the reader is left to wonder whether a model that always predicts the mean value might not actually perform better.
@@
  • Simple Baselines: If a keyword-based baseline performs well, or even better than a proposed method, then the proposed method is likely an overkill. For example, Rouen et al. (2023); Clarkson et al. (2020) showed that even for the prediction of an ESG score (which is a complex target), linguistic features such as the quantity and complexity of text are good predictors. This suggests that simple predictors could perform well even on complex tasks. Indeed, simple baselines can outperform finetuned approaches, for example on TCFD classification (Sampson et al., 2022) or on the ClimaINS dataset (Spokoyny et al., 2023). Beyond a validation (or invalidation) of the proposed approaches, a keyword-based baseline would also allow us to understand if models rely on an “understanding” of the text or merely on word cues. And yet, less than half of the works we studied report the performance of a keyword-based baseline.
@@
  • Human Baselines: Using a human as a baseline can help determine whether a task is well-defined. If even a qualified human cannot achieve high accuracy, this most likely means that the task is subjective, ill-defined, or in disagreement with the dataset. Vice versa, if a human achieves high accuracy, but the model achieves even higher accuracy, this could mean that the model has acquired expert knowledge. However, with the exception of 3 studies, all studies discussed here not come with a human evaluation.
6.2. Model Robustness
@@
Many factors can influence the performance of a model (e.g. data sampling, hyper-parameters, random initialization, model architecture). Thus, a better performance in an experiment could be due also to these factors rather than due to the choice of the model. This is even more true when studies report differences in performance of less than 1%. For this reason, the uncertainty of the evaluation should always be quantified – by confidence intervals, p-values or at least the standard deviation. However, we found that only seven papers, out of the 61 we studied, quantify the uncertainty of the model performance (Webersinke et al., 2022; Leippold and Varini, 2020; Bingler et al., 2023; Schimanski et al., 2024a, 2023b; Lee and Kim, 2023; Schimanski et al., 2023a).
Noisy Text and Adversarial Attacks
Most studies mentioned in this survey trained and evaluated models on datasets that contain only clean examples. However, real-world data is usually noisy. Thus, the models may perform entirely differently on real-world datasets. Studies should quantify not just the performance of the models, but also their robustness, as fine-tuned models are prone to word-based adversarial attacks (Woloszyn et al., 2021).
@@
The class of interest is relatively rare in most tasks. For example, only 5% of text is in the positive class “climate-related” in the study of Yu et al. (2024). This means that (1) a simple baseline that always says “non-climate-related” will have a very high precision, and (2) a model trained on such a dataset would probably overfit. Therefore, most studies mentioned in this review trained and evaluated the models on relatively balanced datasets. On the other hand, a model evaluated only on a balanced dataset may fail when tested on real-world, unbalanced datasets (Friederich et al., 2021). For example, a model reaching an F1-score of 90% on a balanced dataset would significantly drop on an unbalanced dataset with 5% of text in the positive class (to a binary F1-score of 47%, and a macro F1-score of 71%). This is due to the large number of false positives. Thus, studies should consider both the balanced and the unbalanced case.
Precision-Recall Trade-off
In most methods (and most tasks), higher precision can be traded for smaller recall. In the case of greenwashing, both metrics can be useful: a higher recall is desirable when providing leads for human analysis, so as to avoid missing potential greenwashing cases. However, if the goal is to accurately measure greenwashing prevalence, minimizing false positives is critical, as false accusations can have serious repercussions. Most studies, however, are conducted without the use case in mind.
6.3. Data Sources
@@
The studies mentioned in our review use various data sources, most of which are publicly available: annual and sustainable reports  (Doran and Quinn, 2009; Kölbel et al., 2021; Chou et al., 2023; Huang et al., 2020; Kraus et al., 2023; Kheradmand et al., 2021; Rouen et al., 2023; Lai and Chen, 2023; Kang and Kim, 2022; Auzepy et al., 2023; Brié et al., 2022; Webersinke et al., 2022; Luccioni et al., 2020; Kheradmand et al., 2021; Bronzini et al., 2023; Lee and Kim, 2023; Polignano et al., 2022; Kraus et al., 2023), social media (Woloszyn et al., 2021; Luo et al., 2020; Vaid et al., 2022; Coan et al., 2021; Oppong-Tawiah and Webster, 2023; Sharma et al., 2024), specialized websites (Morio and Manning, 2023; Schimanski et al., 2023a; Jin et al., 2022; Leippold and Varini, 2020; Coan et al., 2021). However, multiple works rely on closed-source data: ESG scores, carbon emissions, and machine-readable reports that are collected through private providers such as Refinitiv, Reuters, RavenPack, Trucost, Bloomberg, RobecoSAM, Compustat or Wharton (Sautner et al., 2023; Bingler et al., 2023; Stammbach et al., 2023; Friederich et al., 2021; Brié et al., 2022; Lee and Raschke, 2023; Webersinke et al., 2022; Li et al., 2020; Clarkson et al., 2020; Li et al., 2020; Sautner et al., 2023; Mehra et al., 2022; Sharma et al., 2024; Clarkson et al., 2020; Schimanski et al., 2023b; Vinella et al., 2023; Kraus et al., 2023). Relying on private providers limits the reproducibility of these works, for both academic work and applications.
@@
The most fundamental shortcoming of current work on greenwashing is that there is no comprehensive real-world dataset with instances of greenwashing. This limits the conclusions that can be drawn from current studies drastically. The lack of a gold standard dataset can be explained by several factors:
@@
  • The definitions of greenwashing are often vague and ambiguous, leaving room for interpretation, and making it difficult to annotate examples of greenwashing.
@@
  • Green claims are often not self-contained, but spread across multiple paragraphs in a same document, across multiple documents, or even across different types of sources (such as tables, databases, etc.). Therefore, just collecting all the elements required to correctly identify greenwashing is a challenge on its own.
@@
  • Understanding the claim and annotating greenwashing requires expertise, and sometimes even debate, hence producing greenwashing annotations is a difficult task.
@@
  • Greenwashing is a strong accusation that might impact a company reputation, and that can also, if not substantiated by evidence, be considered a defamation.
To build a dataset of real-world greenwashing cases, several data collection strategies can be considered. To avoid relying on a subjective annotation, researchers can rely on a third-party judgment. For example, using reputable news articles that explicitly discuss greenwashing to find actual examples, drawing on prior research focused on detecting media mentions of greenwashing (Gourier and Mathurin, 2024). Another valuable resource are regulatory records from authorities such as the Advertising Standards Authority (ASA) or databases like greenwash.com.
Missing link to regulatory texts
Existing work often relies on its own definitions of greenwashing. This risks introducing subjectivity and incompatibility with other works. A safer choice is to use regulatory texts as a reference, as this is done, e.g., in Stammbach et al. (2023). Recently, the European Commission adopted the corporate sustainability reporting directive (CSRD) (Union, 2022), making environmental disclosure requirements explicit, and the directive on green claims (US Environmental Protection Agency, 2023), which defines requirements when making green claims. These could be good starting points for defining greenwashing in future research. Such regulations help not just with the definition of greenwashing, but also with its judgment. Instead of emitting a judgment, a method can link a statement to a regulatory text. For example, the works that predict TCFD simply link a statement to a recommendation (Ding et al., 2023; Auzepy et al., 2023; Brié et al., 2022; Bingler et al., 2022; Sampson et al., 2022). This approach makes the methods less prone to ad hoc definitions of scores and judgment strategies, and lets a human take the final decision.
@@
With the exception of Bronzini et al. (2023), the studies reviewed in this work all measure how often a given text contains climate-related content, TCFD disclosures, or specific topics of interest. However, a company can repeatedly emphasize the same objective or repeatedly highlight a single initiative. With this, they skew the quantity of climate-related content in their favor, while they may still fail to implement any meaningful action. This phenomenon has been called “cheap talk” (Bingler et al., 2023). Future research should thus quantify not only how often some pieces of information are conveyed, but also how many different pieces there are.
7. Conclusion
In this survey, we have provided a comprehensive overview of current research initiatives on detection of greenwashing in textual corporate communications. Determining whether a communication contains greenwashing presents three fundamental challenges. First, the absence of a commonly accepted and actionable definition of greenwashing creates ambiguity, complicating any attempt at standardized assessment. Second, there are significant legal and reputational consequences of being accused of greenwashing. Finally, there are no explicit precedents that could serve as jurisprudence for systematic labeling. This makes the annotation of actual examples of greenwashing challenging, and indeed, there is currently no large-scale dataset of instances of greenwashing. For this reason, the detection of greenwashing has remained on a theoretical level without empirical validation (Moodaley and Telukdarie, 2023a). Most existing works target intermediate tasks such as identifying climate-relatedness or green claims. This work also serves its purposes: First, determining whether a text can contain greenwashing at all helps restricting attention to the relevant resources. Second, decomposing greenwashing into more manageable steps can lead to an overall better performance, as supported by the principles behind the Chain-of-Thought (Wei et al., 2022). Third, using frugal approaches helps reduce costs and energy consumption. Finally, intermediate tasks can help humans understand the final verdict, which is important for the explainability of decisions. Several challenges remain in these works:

### v2

@@ 6.1 Evaluation Methodology
One particular issue is that the main hypothesis of the classification task is that categories are distinguishable from each other. This is not always the case. For example, specificity classification [bingler2023cheaptalkspecificitysentiment] is the task of distinguishing specific claims from unspecific claims. It turns out that this task is itself too unspecific, as the low IAA shows. While this issue appears obvious for this particularly ambiguous task, other less ambiguous classes also contain an ambiguous frontier. For example, climate/non-climate labels can be debated in some instance: ”The company produced 10 million computers this year” or ”The company produced 10 million solar panels this year” are respectively predicted ”non-climate” and ”climate” by climateBERT’s climate detection model [bingler2023cheaptalkspecificitysentiment]. In this example, the sentences are fundamentally similar. If the company’s business model is to produce solar panels, then this statement is not necessarily about climate, at least not more than the other one.
@@
Some tasks are not classification tasks, but text generation tasks. Evaluating generated text is a challenging task  [survey_nlg_eval]. We argue that evaluating generated text is still a challenge deserving future research. In the papers we considered for this review, to evaluate generative models, we came across solutions such as using BLEU, METEOR, human-based evaluation, and LLM-based evaluation  [mullappilly-etal-2023-arabic, cliamtebot_2022, s_vaghefi_chatclimate_2023, thulke2024climategpt]. BLEU and METEOR are not great to evaluate tasks such as question answering [chen-etal-2019-evaluating]. Human evaluation also has many shortcomings. Firstly, it is not possible to perform an automatic evaluation; the evaluations are therefore limited to smaller sizes and variations in settings. Secondly, human evaluation of generated texts often relies on comparisons [mullappilly-etal-2023-arabic, thulke2024climategpt]. While thulke2024climategpt refers to specific comparison guidelines [bulian2024assessinglargelanguagemodels], [mullappilly-etal-2023-arabic] relies only on the evaluator’s preferences. Thirdly, human evaluation can be influenced by factors other than the quality of the answer – for example, the observer-expectancy effect [rosenthal1976]. For example, in ChatClimate  [s_vaghefi_chatclimate_2023], humans evaluate the accuracy of text generated by a generic ChatGPT vs. the proposed ChatClimate chatbot. The text produced by the ChatClimate bot always gives specific references (down to page numbers), while the generic ChatGPT does not. It is thus very easy for evaluators to identify the more accurate answer simply by looking at the type of references, so that there is the danger that this was in fact the only facet of the texts that was evaluated (and not, say, the general accuracy or quality of the answer). A third shortcoming of evaluations appears with LLM-based evaluation. Language models are known to reproduce human biases  [gallegos2024biasfairnesslargelanguage]: they are sensitive to the way questions are formulated, to the order of answers in multiple-choice questions  [pezeshkpour-hruschka-2024-large], and to cues that can influence human evaluation such as using “Our Model”/“Their model” (which is the wording used in mullappilly-etal-2023-arabic). LLM-based evaluation should thus always adhere to the same rigorous standards that are applied to human evaluation. Proper control of the evaluation setting is crucial. Subtle factors, such as the specific wording used or the unintentional leakage of task-related information, can bias the results. Therefore, strict protocols and a meticulously designed evaluation environment are essential to mitigate these risks.
Figure 1: Upset plot [upset_plot] of the methods used in the studies mentioned in this literature review. Each bar counts the number of study reporting a given set of methods. The sum of a paper using a given method is reported on the horizontal bars. We have highlighted in red the intersection sets that include both fine-tuned Transformers and keyword-based methods.
@@
Figure 1 shows the approaches that are commonly used in greenwashing detection tasks. We see that the most common solution is fine-tuning transformer models; more recent approaches include using LLMs as few-shot learners. However, when evaluating, the state-of-the-art models are compared to baselines. We argue that some basic baselines are often forgotten, even though they provide important insights:
@@
  • Random/Majority Baseline: Any proposed approach should beat a random baseline. However, we find that only 9% of papers we studied compare their results to the random or majority baseline. For example, Friederich_climate_risk_disclosure reported performances under 50% for a binary classification task, which is worse than a theoretical random classifier on a balanced dataset. Their dataset is not balanced, but it is difficult to draw conclusions without the performance of a random classifier. Another interesting aspect of reporting the random classifier performance is to have a point of reference, or a lower baseline. For example, on SciDCC or ClimaTOPIC, the performance of models is around 50-60%, yet this is excellent performances as the random classifier performed around 5-7% [spokoyny2023answering]. In case of an heavily imbalanced dataset, providing the performance of the majority baseline might be more meaningful. If the dataset is not avaiable, it is not possible to estimate that value, which is the case for [Friederich_climate_risk_disclosure]. Similarly a constant baseline help understand the performance of a regression model. Greenscreen and bronzini_glitter_2023 evaluated the prediction of ESG score, yet they do not provide baseline. If the values are really concentrated, a model which always predict the mean value might have a good performance. Unfortunately, without the distribution of values, it is not possible to estimate it.
@@
  • Simple Baselines: If a keyword-based baseline performs well, or even better than a proposed method, then the proposed method is likely an overkill. And yet, less than half of the works we studied report the performance of a keyword-based baseline. Studies such as rouenEvolutionESGReports2023, clarkson_nlp_us_csr, which showed that linguistic features such as the quantity and complexity of text are good predictors for the ESG score, therefore seemingly simple predictors (e.g. a keyword-based classifier) might perform surprisingly well, even on complex tasks. Moreover, we observed that the simple baseline can even outperform finetuned approaches, for example on TCFD classification [sampson_tcfd-nlp_nodate] or on ClimaINS dataset [spokoyny2023answering]. Beyond a validation (or invalidation) of the proposed approaches, a keyword-based baseline would also allow us to understand if models rely on an “understanding” of the text or merely on word cues.
@@
  • Human Baselines: Using a human as a baseline can help determine whether a task is well-defined. If even a qualified human cannot achieve high accuracy, this most likely means that the task is subjective, ill-defined, or in disagreement with the gold standard. Vice versa, if a human achieves high accuracy, but the model achieves even higher accuracy, this could mean that the model has acquired expert knowledge. However, most of the studies discussed here do not incorporate an evaluation of human performance (Only 3 studies reported Human performances).
6.2 Model Robustness
@@
Many factors can influence the performance of a model (e.g. data sampling, hyper-parameters, random initialization, model architecture). Moreover, we see that many studies report differences in performance of less than 1%. This difference might be due to the choice of model, but it could also be due to the other factors. Quantifying the uncertainty induced by those factors helps understand if a difference is significant. This could be done using confidence intervals, p-values or simply reporting the standard deviation. However, we found only seven papers that quantify the uncertainty of the model performance  [nicolas_webersinke_climatebert_2021, varini_climatext_2020, bingler2023cheaptalkspecificitysentiment, Schimanski2024nature, schimanski_bridging_2023, LEE2023119726, tobias_schimanski_climatebert-netzero_2023].
Adversarial Attacks and Noisy Text
While fine-tuning leads to the best performances, it is also prone to word-based adversarial attacks  [vinicius_woloszyn_towards_2021], as the training data are often relatively small, and might not include syntactically diverse examples. Noisy texts are an inherent part of real-world applications. Yet, most studies mentioned in this review trained and evaluated models on clean datasets (containing only clean examples) and did not quantify the robustness of their approaches to noise.
@@
In most of the tasks, the class of interest is relatively rare. For example, only 5% of text is in the positive class “climate-related” in the study of yu_climatebug_2024. This means that (1) a simple baseline that always says “non-climate-related” will have a very high precision, and (2) a model trained on such a dataset would probably overfit. Therefore, most studies mentioned in this review trained and evaluated the models on relatively balanced datasets. On the contrary, a model that was evaluated only on a balanced dataset may fail spectacularly when tested on real-world, unbalanced datasets  [Friederich_climate_risk_disclosure]. For example, a model reaching an F1-score of 90% on a balanced dataset would significantly drop on an unbalanced dataset with 5% of text in the positive class (to a binary F1-score of 47%, and a macro F1-score of 71%). This is due to the large number of false positives. For tasks related to greenwashing detection, annotating data is a time-consuming and resource-intensive process, often resulting in an imbalanced dataset that fails to represent all relevant cases adequately. This imbalance complicates the model’s performance trade-off: a higher recall is desirable when providing leads for human analysis to avoid missing potential greenwashing cases. However, if the goal is to accurately measure greenwashing prevalence, minimizing false positives is critical, as false accusations can have serious repercussions.
6.3 Data Sources
@@
The studies mentioned in our review use various data sources, most of which are publicly available (annual and sustainable reports used in 18 studies)social media (6 studies)specialized website (5 studies). However, multiple works rely on closed-source data, this include ESG score, carbon emissions, machine-readable reports that are collected through private providers such as Refinitiv, Reuters, RavenPack, Trucost, Bloomberg, RobecoSAM, Compustat or Wharton (17 studies). Relying on private providers limits the reproducibility of those works, for both academic work and applications.
@@
While theoretical frameworks have provided a starting point for defining greenwashing, a critical gap remains: the lack of a comprehensive, real-world dataset. This can be explained by several factors:
@@
  • Greenwashing definitions are often vague and ambiguous, leaving room for interpretation, making it difficult to use for annotating examples of greenwashing
@@
  • Green claims are often not self-contained, meaning that the full content of information required to judge if a claim is misleading or not is spread: across multiple paragraphs in a same document, or even across multiple documents. Therefore collecting all the elements required to correctly identify greenwashing is a challenges on its own.
@@
  • Understanding the claim and annotating greenwashing requires expertise, and sometimes debates to make sure the communication is misleading, making the annotation task even harder.
@@
  • Greenwashing is a strong accusation, that might impact a company reputation, and that can also, if not substantiated with evidence, be considered as defamation.
To address these challenges, several potential data collection strategies can be considered. To avoid relying on a subjective annotation, researchers can rely on a third-party judgement. For example, using news articles that explicitly discuss greenwashing, drawing on prior research focused on detecting media mentions of greenwashing  [gourierGreenwashingIndex2024]. Another valuable resource could include regulatory records from authorities such as the Advertising Standards Authority (ASA) or databases like greenwash.com. Alternatively, previous works tried to produce signals that could potentially indicate greenwashing, future research might rely on such indicators to first filter-out texts that are potentially containing greenwashing, and annotate those selected texts. Intermediary tasks are also important for other reasons. First, decomposing tasks often results in better performance, as supported by the principles behind Chain-of-Thought. Second, when AI systems are designed to include human oversight, it is crucial for these intermediary tasks to remain understandable and accessible to users. Thirdly, the design of energy-efficient AI systems can benefit from employing simpler models for tasks where high precision is not required, such as document retrieval for RAG. Finally, relying on intermediary tasks for the evaluation of models’ performances is necessary for both understanding better the models, their limits, and to propose auditable pipelines with humans in the loop.
Relying on regulatory texts
As mentioned above, two of the challenges in building a dataset of examples of greenwashing are using definitions that are subjective and mitigating risks of defamation. Using regulatory texts as reference helps mitigate those two issues. Firstly, it can help build a proper definition of the label. For example, stammbach_environmental_2023 rely on the definition of green claims by the European Commission to build their definition of environmental claims. Secondly, instead of emitting a judgment, the method can link a statement to a regulatory text. For example, all works focusing on TCFD simply link a statement to a recommendation [dingCarbonEmissionsTCFD2023, auzepy_evaluating_2023, bjarne_brie_mandatory_2022, bingler_cheap_2021, sampson_tcfd-nlp_nodate]. Recently, the European Commission adopted the CSRD which defines more precisely greenwashing. Future research could start by finding content related to parts of the regulatory texts related to greenwashing.
@@
With the exception of bronzini_glitter_2023, the studies reviewed in this work primarily focus on quantifying textual content related to climate, TCFD disclosures, or specific topics of interest. However, as highlighted by bingler2023cheaptalkspecificitysentiment, companies increasingly engage in ”cheap talk,” where statements lack substantive commitments or actionable details. This underscores the need to assess not only the volume of climate-related communication but also the informational content conveyed. For instance, a company might repeatedly emphasize its objectives or highlight a single initiative, creating the illusion of widespread organizational change, while failing to implement meaningful action. Future research should focus on quantifying the information instead of the amount of text related to a topic.
Restricting the definition of greenwashing
Except for obvious lies, what counts as “misleading communication” is intrinsically subjective. For some, disproportionately highlighting a company’s green initiatives is misleading (LEE_greenwashing); others interpret the same behavior as employees’ pride in early achievements or as a strategic shift to engage stakeholders around more ambitious climate action. Likewise, hyewon_kang_analyzing_2022 frame greenwashing as overly positive texts, yet such positivity can be a stylistic choice rather than deceptive intent. Moving from a broad to a narrower definition can improve the objectivity of what is considered greenwashing. The trade-off with narrower definitions is twofold: (i) they may exclude other forms of greenwashing, and (ii) the definition itself can then be attacked for not adequately defining greenwashing. Still, this is preferable to relying on a broad but vague definition, which would produce unreliable and subjective annotations. Future work should further explore these and other narrow but practical definitions, as they remain useful starting points for uncovering greenwashing practices.
7 Conclusion
In this survey, we have provided a comprehensive overview of current research initiatives on detection of greenwashing in textual corporate communications. Determining whether a communication contains greenwashing presents three fundamental challenges. First, the absence of a commonly accepted and actionable definition of greenwashing creates ambiguity, complicating any attempt at standardized assessment. Second, there are significant legal and reputational consequences of being accused of greenwashing. Finally, there are no explicit precedents that could serve as jurisprudence for systematic labeling. This makes the annotation of actual examples of greenwashing challenging, and indeed, there is currently no large-scale dataset of instances of greenwashing. For this reason, the detection of greenwashing has remained on a theoretical level without empirical validation [moodaley_conceptual_2023]. Most existing works target intermediate tasks such as identifying climate-relatedness or green claims. This work also serves its purposes: First, determining whether a text can contain greenwashing at all helps restricting attention to the relevant resources. Second, decomposing greenwashing into more manageable steps can lead to an overall better performance, as supported by the principles behind the Chain-of-Thought [cot]. Third, using frugal approaches helps reduce costs and energy consumption. Finally, intermediate tasks can help humans understand the final verdict, which is important for the explainability of decisions. Several challenges remain in these works: